OpenStudy (anonymous):

Use the matrix method to solve: x + 4y = 8 2x + y = 9 x=4 y=?

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmate):

You can solve it by Gaussian elimination, or by Cramer's method. Which one have you done at school?

OpenStudy (anonymous):

none

OpenStudy (anonymous):

x is 4 idk y

OpenStudy (mathmate):

Well, since you already know x, you don't have to use matrix method to solve for y. Put x=4 in one of the two equations to reduce the number of unknowns to one, solve for y.

OpenStudy (anonymous):

i really dont know how to this the last question

OpenStudy (mathmate):

Given x+4y=8 and you know x=4, so put x=4 in the above equation to get 4+4y=8 and try to solve for y.

OpenStudy (anonymous):

ok so 4 cancels

OpenStudy (mathmate):

Not really, unless you mean you would subtract 4 from both sides.

OpenStudy (anonymous):

yeah from both sides

OpenStudy (mathmate):

So what do you get for y?

OpenStudy (anonymous):

8

OpenStudy (mathmate):

Starting with 4+4y=8 Subtract 4 from each side: 4-4 +4y = 8-4 4y = 4 Can you now solve for y?